OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y =1/5 x - 10. y = 5x + 10 y = 5x y = -5x - 10 y = -5x

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

An equation of the line : \[y=mx+b\] And because this line is perpendicular to y=1/5x-10 ,then : \[-1=\frac15\times m\] And then : \[m=-5\] Now to find be we have : \[-5=m\times 1+b\] So : \[b=-5-m=-5-(-5)=-5+5=0\] So an equation of the line is : \[y=-5x\]

OpenStudy (anonymous):

Thanks :) Help me with some more please ?

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

I am here !

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2 y - 4 =1/2 (x - 1) y + 4 = 2(x - 1) y + 4 =1/2 (x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

The slope is : \[m=\frac12\] So : \[\frac{y-(-4)}{x-1}=m=\frac12\] So : \[y+4=\frac12(x-1)\]

OpenStudy (anonymous):

Thanks :) Theres only a few more

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

An equation of the line is : \[y=mx+b\] Since the line is parallel to y=-x-4 then m=-1 To find b we have : \[1=m\times (-3)+b\] Then : \[b=1+3m=1+3\times(-1)=1-3=-2\] An equation is : \[y=-x-2\]

OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

It is like the previous one

OpenStudy (anonymous):

Choose the equation of the vertical line passing through the point (8, -2). x = -2 y = -2 x = 8 y = 8

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

Since the line is vertical then : \[x=b\] To find b, since (8,-2) is from the line then \[8=b\]

OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21 @Noura11

OpenStudy (anonymous):

To find the x-intercept we solve : \[x=0\] this means : \[-21-7y=0\] And : \[y=\frac{-21}{7}=-3\] So, the x-intercept is the point (0,-3) Now : To find the y-intercept we solve : \[y=0\] this means : \[-\frac17x-3=0\] And : \[x=-21\] So, the y-intercept is the point (-21,0)

OpenStudy (anonymous):

wait, -21 is the x or y intercept ? @Noura11

OpenStudy (anonymous):

The x-intercept is : (-21,0) and the y-intercept is (0,-3) I have done a mistake ! :)

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2. y =1/2 x + 5 y =1/2 x - 3 y =1/2 x - 5 y =1/2 x + 3

OpenStudy (anonymous):

@Noura11

OpenStudy (anonymous):

Is it not clear to you ?

OpenStudy (anonymous):

I dont understand -.- @Noura11

OpenStudy (anonymous):

You have to try to solve it, even it is false !

OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)? @Noura11

OpenStudy (anonymous):

The slope m is calculated by : \[m=\frac{y_2-y_1}{x_2-x_1}\]